NONLINEAR SPECTROSCOPIC PROBES OF UNIMOLECULAR DYNAMICS: VIBRATIONALLY-MEDIATED TUNNELING IN X̃^{1}A_{1} TROPOLONE

\underline{D. MURDOCK}, L. A. BURNS, and P. H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520-8107.

The mode specificity of proton-transfer dynamics in the ground state (X̃^{1}A_{1}) of tropolone has been explored at nearrotational resolution by implementing a fully coherent variant of stimulated emission pumping within the framework of two-color resonant four-wave mixing spectroscopy (TC-RFWM). Ongoing studies have exploited rovibronically-resolved features of the Ã−X̃ origin band as a “doorway” for selectively interrogating vibrationally-excited levels of the ground electronic manifold, with judicious selection of incident/detected polarization characteristics affording a means for discriminating rotational branches and alleviating spectral congestion.^{a} Several mid-range vibrational transitions (E_{vib} ≤ 1700 cm^{−1}) have been interrogated under ambient, bulk-gas conditions, with term energies determined for the symmetric and antisymmetric (tunneling) components enabling the attendant tunneling-induced bifurcations to be extracted.^{b} The dependence of tunneling rate (or hydron migration efficiency) on vibrational motion is discussed in terms of corresponding atomic displacements and permutation-inversion symmetries for the tropolone skeleton.

aA. E. Bracamonte and P. H. Vaccaro, J. Chem. Phys. 119(2), 887 (2003); 120(10), 4638 (2004).

bD. Murdock, L. A. Burns and P. H. Vaccaro, J. Chem. Phys. 127, 081101 (2007).